OpenStudy (anonymous):

Trig help? Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0

OpenStudy (anonymous):

cos 2x = 2sinxcosx By factorizing cos x out: cos 2x - cos x = 0 2sinxcosx - cosx = 0 (cosx)(2sinx-1) = 0 Solve for: cosx = 0 AND 2sinx -1 = 0 x=??? AND sinx = 1/2 x=??? AND x2=??? Find which are the common solutions, that will be your answer.

OpenStudy (anonymous):

thats wrong... you have to use cos2x=2(cosx)^2-1 sin2x=2sinxcosx....not cos2x

OpenStudy (anonymous):

yeah o.o

OpenStudy (anonymous):